Ka and Kb are inversely related. Thus, if Ka is small, Kb must be large. However, the textbook says that (CO3)2- is a weak base and its conjugate, HCO3-, is a weak acid. How is this possible? Since HCO3- is a weak acid, Ka should be small. Since Ka and Kb are inversely related, shouldnt the Kb of the conjugate base to the weak acid HCO3- be large, making it a strong base?
Answer #1

Higher the Ka or Kb means higher strength of acid or base.

Ka * Kb = Kw = 1*10^-14

Ka = (1*10^-14) / Kb

Hence, Ka 1/Kb

Ka is of an acid and Kb is of its conjugate base.

HCO3- <-----> CO3-- + H+

Acid Conjugate base

Ka of HCO3- = 4.8*10^-11 that means HCO3- dissocation is very less.

Ka = [CO3--][H+] / [HCO3-]

[CO3--][H+] / [HCO3-] = 4.8*10^-11

[CO3--][H+] = ~ 4.8*10^-11 almost

Thus H+ ion concentration is very low and that makes HCO3- a weak acid.

Kb of CO3-- (please note that here we should not consider Kb of HCO3-, instead we should consider Kb of CO3--)

Kb of CO3-- = 1.8*10^-4

So, comparatively Ka of HCO3- <<< Kb of CO3--.

However CO3-- is still a weaker base because

CO3-- + H2O <-----> HCO3- + OH-

does not make much of OH- ion concentration increase.

Thus the relationship given in the book is correct.